What is the equation for the following graph?
parabola opening to the right, with vertex at (1, 2) and directrix at x equals negative 1

Answer by lwsshak3(11628)

What is the equation for the following graph?
parabola opening to the right, with vertex at (1, 2) and directrix at x equals negative 1
..
Standard form for parabola opening to the right:
(y-k)^2=4p(x-h),with (h,k) being the (x,y) coordinates of the vertex which are given.
axis of symmetry,y=2
On the axis of symmetry,p=1/2 the distance between the vertex and directrix.
Distance between vertex and directrix=2
Therefore, p=1
Equation:
(y-2)^2=4(x-1)
See the graph below:
..
y=2+(4(x-1))^.5